### v1

# Shadow of Quantum Improved Regular Kerr Black Hole and parameter constrains with EHT observations

Abstract
Quantum Improved Regular Kerr (QIRK) Black Hole is a rotating regular black hole based on the asymptotic safety method. This black hole not only resolves ring singularity and avoids closed timelike curves, but also has well defined thermodynamics. Therefore, it is crucial to find some observable features of this rotating black hole. In this article, we numerically determine the specific parameter range of the QIRK black hole after ensuring the three key properties mentioned above, while investigating its black hole shadow, we find that the extremal QIRK black hole, under a critical angular momentum $a_{\mathrm{cri}}$, can have a similar shadow to the non extremal Kerr black hole. Furthermore, with recent observations from the Event Horizon Telescope (EHT) of Sgr A* and earlier observations of the supermassive black hole M87*, we constrain the QIRK black hole using observational data and explore its potential as an astronomical object.
preprint: USTC-ICTS/PCFT-24-39
Xia-Yuan Liub444e-mail address: liuxiayuan@mail.ustc.edu.cn
October 21, 2024
I Introduction
Black holes, predicted by general relativity, are one of the most enigmatic objects in the universe. Classical black holes, such as the Kerr black hole kerr_gravitational_1963 which describes a rotating black hole, play a crucial role in understanding numerous astrophysical phenomena, including quasars and gamma-ray bursts. However, a significant issue with these classical solutions is the presence of singularities, characterized by the divergence of curvature invariants and geodesic incompleteness wald_general_1984 ; hawking_large_1973 . To address this problem, regular black holes (RBHs) were introduced. Unlike traditional black holes, they are free from essential singularities throughout the entire spacetime bardeen_non-singular_1968 ; dymnikova_vacuum_1992 ; ayon-beato_regular_1998 ; bronnikov_regular_2001 . The first regular black hole model was proposed by Bardeen bardeen_non-singular_1968 , and was later interpreted as a solution derived from Einstein’s field equations with nonlinear electromagnetic field Ayon-Beato:2000mjt . This approach has been extended to explain various spherically symmetric RBH models Ayon-Beato:2000mjt ; PhysRevLett.96.031103 ; bronnikov_regular_2001 ; fan_construction_2016 ; bronnikov_regular_2006 ; bronnikov_field_2022 . Since Bardeen’s work, numerous other regular black hole models have been developed Ayon-Beato:1999kuh ; burinskii_new_2002 ; dymnikova_regular_2004 ; berej_regular_2006 ; junior_regular_2015 ; sajadi_nonlinear_2017 ; toshmatov_comment_2018 ; bambi_rotating_2013 ; toshmatov_rotating_2014 ; azreg-ainou_generating_2014 ; ghosh_nonsingular_2015 ; ghosh_radiating_2015 , including several rotating RBH models bambi_rotating_2013 ; toshmatov_rotating_2014 ; azreg-ainou_generating_2014 ; ghosh_nonsingular_2015 ; ghosh_radiating_2015 .
Two primary methods have been established for constructing regular black hole models. The first one involves solving Einstein’s field equations with specific matter sources noauthor_noncommutative_2006 ; noauthor_non-commutative_2009 ; noauthor_non-minimal_2007 . The second method derives RBHs as quantum corrections to classical black holes with singularities. The notable examples come from frameworks such as loop quantum gravitymodesto_disappearance_2004 ; gambini_black_2008 ; perez_black_2017 . A promising alternative is the asymptotic safety scenario Reuter:1996cp ; Souma:1999at ; doi:10.1142/10369 ; 2019qgfr.book…..R ; bonanno_renormalization_2000 ; koch_black_2014 ; chen2024quantumimprovedregularkerr , which suggests that quantum gravity effects, particularly through the functional renormalization group, introduce a repulsive force near black hole cores bonanno_renormalization_2000 ; Bonanno:2006eu , potentially resolving singularities. Many such approaches have been studied for various black hole solutions reuter_running_2004 ; falls_black_2012 ; platania_black_2023 ; eichhorn_black_2022 ; koch_structural_2013 ; pawlowski_quantum-improved_2018 ; ishibashi_quantum_2021 ; Litim:2013gga ; Harst:2011zx ; Falls:2012nd ; Adeifeoba:2018ydh ; 10.3389/fphy.2020.00188 ; Borissova:2022mgd ; Ruiz:2021qfp .
The Quantum Improved Regular Kerr (QIRK) black hole model is based on the asymptotic safety framework, introducing a running gravitational coupling $G(r)$ that varies with energy scale 10.3389/fphy.2020.00056 . This approach not only resolves the Kerr black hole’s ring singularity but also addresses issues such as closed timelike curves (CTCs) carter_global_1968 , which would otherwise violate causality. Unlike classical black holes, which suffer from infinite curvature at singularities, the quantum-improved model ensures finite curvature invariants while maintaining consistent thermodynamic properties PhysRevD.83.044041 ; chen_running_2022 . This is crucial, as it allows the black hole to obey the fundamental laws of thermodynamics even as quantum gravitational corrections stabilize its core.
The observation of light deflection in a gravitational field in 1919 provided the first experimental confirmation of a prediction from general relativity. This phenomenon, now known as gravitational lensing, is a key method for detecting weak gravitational fields. When a photon passes near a black hole, it can either be absorbed, bounded in its vicinity, or scattered to infinity. The observed image of such photons is referred to as the black hole shadow. The first analytic study of black hole shadows began with Synge’s Synge:1966okc and Luminet’s Luminet:1979nyg discussions of the Schwarzschild black hole, followed by Bardeen’s analysis of Kerr black hole in 1973 1973blho.conf..215B . These foundational works paved the way for further exploration of black hole shadows in various models noauthor_escape_nodate ; grenzebach_photon_2014 ; bambi_apparent_2009 ; takahashi_shapes_2004 ; wei_observing_2013 ; abdujabbarov_shadow_2013 ; atamurotov_optical_2015 ; schee_optical_2009 ; tsukamoto_constraining_2014 ; takahashi_black_2005 ; bambi_direct_2012 ; khodadi_black_2020 ; shaikh_shadows_2019 ; falcke_toward_2013 ; tsupko_analytical_2017 ; qian_cuspy_2022 ; noauthor_charged_2022 ; contreras_black_2020 ; perlick_calculating_2022 ; falcke_viewing_1999 ; atamurotov_shadow_2013 ; vries_apparent_2000 ; hioki_measurement_2009 ; abdujabbarov_coordinate-independent_2015 ; yumoto_shadows_2012 ; Cunha:2015yba ; Cunha:2018acu ; Lima:2021las ; Chen:2020aix ; 2005Natur.438…62S ; Cao:2024kht , including regular black holes li_measuring_2014 ; amir_shapes_2016 ; kumar_shadow_2019 ; banerjee_signatures_2022 ; allahyari_magnetically_2020 ; olmo_shadows_2023 ; abdujabbarov_shadow_2016 ; tsukamoto_black_2018 ; dymnikova_identification_2019 ; noauthor_ergosphere_2020 ; jusufi_quasinormal_2021 ; ling_shadows_2022 ; Kumar:2018ple ; Kumar:2020yem ; Kumar:2020ltt ; Cao:2023par .
A major breakthrough in the study of black hole shadows came in 2019 EventHorizonTelescope:2019dse ; EventHorizonTelescope:2019ggy ; EventHorizonTelescope:2019pgp ; EventHorizonTelescope:2019ths ; EventHorizonTelescope:2019uob ; EventHorizonTelescope:2019jan when the Event Horizon Telescope (EHT) released the first image of the supermassive black hole M87*. This was followed in 2022 by EHT’s observations of Sgr A* EventHorizonTelescope:2022wkp ; EventHorizonTelescope:2022xqj ; EventHorizonTelescope:2022urf ; EventHorizonTelescope:2022wok ; EventHorizonTelescope:2022apq ; EventHorizonTelescope:2022exc . Based on these EHT results, the shadows of M87* and Sgr A* have become important tools for testing and constraining theories of gravity in the strong-field regime. Recently, a vast body of literature aimed at deepening our understanding of black holes through these observations afrin_parameter_2021 ; vagnozzi_hunting_2019 ; afrin_estimating_2022 ; vagnozzi_horizon-scale_2023 ; kumar_rotating_2020 ; kumar_gravitational_2020 ; walia_testing_2022 ; ghosh_parameters_2021 ; islam_investigating_2023 ; noauthor_shadows_nodate ; walia_observational_2023 ; Cunha:2019ikd ; Kumar:2020yem ; Kumar:2020ltt ; Uniyal:2022vdu ; Pantig:2022ely ; Afrin:2021wlj ; Bambi:2019tjh .
Given the importance of black hole shadows and the recent EHT observations, this article investigates the QIRK black hole. By hypothesizing that real black holes may conform to this model, we explore its expected observational signatures. Our research highlights the novel properties of the QIRK black hole’s shadow, particularly under extremal conditions, where it significantly differs from that of an extremal Kerr black hole. Furthermore, we use EHT data to constrain the parameters of QIRK black holes, aiming for a more comprehensive and detailed understanding of this model.
This article is organized as follows: Section II examines the constraints imposed by QIRK black holes, including their horizons, regularity, and avoidance of closed timelike curves. Section III discusses the motion of light in both the radial and angular directions. Section IV introduces the shadow correlation function, which facilitates the discussion of shadows in Section V. Section VI analyzes black hole shadows by using EHT data to further constrain the parameters of QIRK black holes. Finally, Section VII summarizes the article and discusses the potential candidature of QIRK black holes, through shadow observations, as viable models for astrophysical black holes.
II quantum improved Regular Kerr Black Hole
The asymptotically safe scenario for a quantum generalization of general relativity with a cosmological constant proposes an energy-scale $k$ dependent Newton constant $G(k)$ and cosmological constant $\Lambda(k)$. Assuming the cosmological constant is already at its fixed point and negligibly small, the solution for the Newtonian coupling derived from the associated renormalization group equations takes the form pawlowski_quantum-improved_2018 :
$G(k)=\frac{G_{0}}{1+\omega G_{0}k^{2}},$ (1)
where $\omega$ is a constant of order 1 representing quantum effects, and $G_{0}=G(0)$ is the Newton constant.
Based on resolving singularities and ensuring the consistency of black hole thermodynamics, Ref. chen2024quantumimprovedregularkerr also incorporates some physical discussions, namely that quantum corrections in graviy theory should become significant at the Planck scale. They adopted the following identification:
$k^{2}=\frac{\xi^{2}}{(r^{2}+a^{2})(Mr)^{p}},$ (2)
where $\xi$ is a dimensionless parameter. By setting $\widetilde{\omega}=\xi^{2}\omega$, we obtain a specific form for the Newton coupling $G(r)$, which allows us to construct the QIRK metric in Boyer-Lindquist coordinates chen2024quantumimprovedregularkerr
$\mathrm{d}s^{2}=-\left(1-\frac{2G(r)Mr}{\Sigma}\right)\mathrm{d}t^{2}+\frac{\Sigma}{\Delta}\mathrm{d}r^{2}-\frac{4aG(r)Mr\sin^{2}\theta}{\Sigma}\mathrm{d}t\mathrm{d}\varphi+\Sigma~{}\mathrm{d}\theta^{2}+\frac{((r^{2}+a^{2})^{2}-a^{2}\Delta\sin^{2}\theta)\sin^{2}\theta}{\Sigma}\mathrm{d}\varphi^{2}\,,$ (3)
where
$\Sigma=r^{2}+a^{2}\cos^{2}\theta\,,\qquad\Delta=r^{2}-2G(r)Mr+a^{2}\,,$ (4)
and
$G(r)=G_{0}\frac{(Mr)^{p}(r^{2}+a^{2})}{(r^{2}+a^{2})(Mr)^{p}+\widetilde{\omega}G_{0}}\,.$ (5)
Here, $G_{0}$ is the Newton constant, $M$ is the mass parameter, $a$ is the rotation parameter, and $\widetilde{\omega}$ is the quantum correction parameter introduced by the QIRK metric. The magnitude of $\widetilde{\omega}$ reflects the strength of the quantum effects, and we require $\widetilde{\omega}\geq 0$. When $\widetilde{\omega}=0$, the solution reduces to the classical Kerr case.
For ease of discussion, we will introduce dimensionless physical quantities as follows:
$x=\frac{r}{G_{0}M},\quad A=\frac{a}{G_{0}M},\quad\widetilde{\Omega}=\frac{\widetilde{\omega}}{G_{0}^{1+p}M^{2+2p}}.$ (6)
With these substitutions, the function $\Delta(r)$ becomes
$\frac{\Delta(r)}{(G_{0}M)^{2}}=x^{2}-2\frac{x^{p}(x^{2}+A^{2})}{x^{p}(x^{2}+A^{2})+\widetilde{\Omega}}x+A^{2}.$ (7)
In the following discussion, we will use the dimensionless variables $x$, $A$, and $\widetilde{\Omega}$. However, for convenience, we will still denote them by the original symbols $r$, $a$, and $\widetilde{\omega}$.
In this section, we will discuss the specific conditions required to resolve the singularities, ensure the existence of horizons and avoid the closed timelike curves. We will examine the implications of each condition and their corresponding physical significance. Finally, these elements will be integrated to provide a comprehensive understanding of the QIRK metric.
II.1 Resolving the Singularity and Ensuring Horizon Existence
As a regular black hole, one of the key requirements is ensuring that the spacetime remains regular, even at the center, where classical singularity would typically emerge. To understand how this regularity is maintained, we focus on resolving the singularity by examining the behavior of the metric and curvature invariants near $r=0$.
For the rotating black holes, the algebraically complete set consists of four invariants, two of them, $R$ and $I_{6}$, are real, while the other two, $I$ and $K$, are complex chen2024quantumimprovedregularkerr ; torres_regular_2016 . The resolution of singularity has been discussed in detail in Ref. chen2024quantumimprovedregularkerr , thus we will only briefly explain it here. The behavior of the four curvature invariants near the center of the equatorial plane from different directions , i.e. $r\rightarrow 0$, have the following fall off:
$R\sim\left\{\begin{array}[]{ll}r^{p-1},&\theta\neq\frac{\pi}{2}\\ r^{p-3},&\theta=\frac{\pi}{2}\end{array}\right.,\quad I_{6}\sim I\sim\left\{\begin{array}[]{ll}r^{2(p-1)},&\theta\neq\frac{\pi}{2}\\ r^{2(p-3)},&\theta=\frac{\pi}{2}\end{array}\right.,\quad K\sim\left\{\begin{array}[]{ll}r^{3(p-1)},&\theta\neq\frac{\pi}{2}\\ r^{3(p-3)},&\theta=\frac{\pi}{2}\end{array}\right..$
In order to avoid singularities and discontinuities, it is necessary to set $p$ as an integer greater than 3. Furthermore, if $p$ is an odd integer, the term $(r^{2}+a^{2})(Mr)^{p}+\omega G_{0}$, which appears in the denominator of all four invariants (see Eqs. (4.8) - (4.11) in Ref. chen2024quantumimprovedregularkerr ), will have a zero point when $r<0$ (both $M$ and $\omega$ are positive). At this zero point, these invariants will diverge, resulting in a curvature singularity. Therefore, we consider the case where $p$ is an even integer greater than 3.
If we disregard the specific form of $G(r)$ and examine the behavior of the four invariants (see Eqs. (A5)-(A9) in chen2024quantumimprovedregularkerr ) as $r\rightarrow 0$, we obtain the following conditions:
$G(0)=0,~{}~{}~{}~{}~{}G^{\prime}(0)=0,~{}~{}~{}~{}~{}G^{\prime\prime}(0)=0,~{}~{}~{}~{}~{}G^{\prime\prime\prime}(0)=0.$ (8)
These constraints are necessary to resolve the singularity.
Since we are interested in the QIRK black hole rather than a compact star without a horizon, our primary objective is to identify the parameter range within which horizons exist. The horizons are determined by the equation
$\Delta=r^{2}-2G(r)r+a^{2}=0\,.$ (9)
Fig. 1 illustrates the relationship between $\Delta(r)$ and $r$. As $r\rightarrow\pm\infty$, $\Delta(r)$ approaches +$\infty$+ ∞, so only a limited range of $r$ values is shown. From Fig. 1, we observe that this black hole exhibits an additional zero point in $\Delta^{\prime}(r)$ compared to the Kerr black hole with the corresponding rotation parameter $a$.
For fixed values of $a$ and $p$, Fig. 1 illustrates that $\widetilde{\omega}$ reaches a critical threshold, $\widetilde{\omega}_{c}$, at which the inner and outer horizons coincide. This represents the extremal case. If $\widetilde{\omega}$ exceeds $\widetilde{\omega}_{c}$, the metric describes a compact object without a horizon. The parameter $\widetilde{\omega}$ reflects the strength of quantum corrections. Thus, the existence of a horizon requires moderate quantum effects, specifically $\widetilde{\omega}\leq\widetilde{\omega}_{c}$. When this condition is met, the metric corresponds to a black hole, which is the primary focus here. It is also worth noting that Fig. 1 represents the particular case of $a=0.9$ and $p=4$. As $a$ varies, (9) shows that the functions in Fig. 1 undergo an overall vertical shift. Therefore, similar to the Kerr black hole, the case where $a>1$ is excluded. Changes in $p$ exhibit behavior analogous to that shown in Fig. 1.
Figure 1: we’ve chosed $a=0.9$ and $p=4$ to analysize the effect of $\widetilde{\omega}$ on the behavior of $\Delta(r)$. $\widetilde{\omega}_{c}$ is the critical vaule which two horizons will coincident. If $\widetilde{\omega}>\widetilde{\omega}_{c}$, then there are no horizons. On the contrary, there exists at least one horizon.
II.2 Avoiding Closed Timelike Curves
It is well known that closed timelike curves (CTCs) exist in the Kerr black hole, which could lead to violations of causality. However, the CTCs are eliminated in this black hole model due to quantum effects. In this section, we examine the parameter constraints that arise from the avoidance of CTCs.
The issue of avoiding CTCs was also addressed in Ref. chen2024quantumimprovedregularkerr . However, due to the $p$-th power in the function $G(r)$, analytical calculations become very complicated. Therefore, they applied certain approximations to derive a sufficient condition for avoiding CTCs:
$1\displaystyle 1<\frac{\widetilde{\omega}^{1/p}}{G_{0}M^{2}},\qquad\frac{a^{2}}{G_{0}}<$ 1 . (10)
Firstly, considering that the $g_{\varphi\varphi}$ component contains the $\Delta(r)$ term, even when $\theta=\pi/2$, its expression remains challenging to simplify due to the higher powers of $G(r)$ in $\Delta(r)$. Secondly, to accurately determine the parameter ranges that allow us to utilize the Event Horizon Telescope (EHT) data for tighter parameter constraints in Sec. VI, we adopt a numerical approach to analyze the conditions required to avoid CTCs.
Specifically, we consider the condition $g_{\varphi\varphi}>0$ to avoid CTCs:
$g_{\varphi\varphi}=\frac{((r^{2}+a^{2})^{2}-a^{2}\Delta\sin^{2}\theta)\sin^{2}\theta}{\Sigma},\qquad\Sigma=r^{2}+a^{2}\cos^{2}\theta.$ (11)
A sufficient condition can be obtained by setting $\theta=\pi/2$. Since $G(r)$ is always positive, we only need to consider the case where $r<0$. In this case, Eq. (11) reduces to:
$g_{\varphi\varphi}=\frac{r^{3}+a^{2}r+2a^{2}G(r)}{r}\sim\left\{\begin{array}[]{ll}+\infty,&r\rightarrow-\infty,\\ a^{2},&r\rightarrow 0.\end{array}\right.$ (14)
Next, let us consider the polynomial $\psi(r)=r^{3}+a^{2}r+2a^{2}G(r)$. To ensure $g_{\varphi\varphi}>0$ for $r<0$, $\psi(r)$ must have no zero points. Fig. 2 plots $\psi(r)$ for $a=0.9$ and $p=4$ with varying $\widetilde{\omega}$ values. As shown, a minimum value of $\widetilde{\omega}$, denoted $\widetilde{\omega}_{m}$, exists such that $\psi(r)$ is tangent to the $r$-axis. Thus, to avoid CTCs, $\widetilde{\omega}$ must satisfy $\widetilde{\omega}>\widetilde{\omega}_{m}$.
Therefore, in order to avoid CTCs, a lower bound on the quantum correction was proposed, specifically requiring $\widetilde{\omega}>\widetilde{\omega}_{m}$. Additionally, $\widetilde{\omega}_{m}$ exists for other values of $a$ and $p$, similar to the previous discussion regarding $\widetilde{\omega}_{c}$.
Figure 2: Plot of $\psi(r)$ as a function of $r$ for $a=0.9$ and $p=4$. When $\widetilde{\omega}>\widetilde{\omega}_{m}$, $\psi(r)$ has no zero points, corresponding to $g_{\varphi\varphi}>0$ for all $r$. Conversely, if $\psi(r)$ has a zero point, there will exist some $r$ for which $g_{\varphi\varphi}<0$, implying the existence of CTCs.
While ensuring that the above three conditions are satisfied, we obtain the parameter range illustrated in Fig. 3. The black curve represents $\widetilde{\omega}=\widetilde{\omega}_{c}$, where the horizons coincide, and the orange curve corresponds to $\widetilde{\omega}=\widetilde{\omega}_{m}$, where $\psi(r)$ becomes tangent to the $r$-axis. The region of interest is where $\widetilde{\omega}$ lies between $\widetilde{\omega}_{c}$ and $\widetilde{\omega}_{m}$, as indicated by the shaded region in Fig. 3.
Additionally, the green-shaded area denotes the region where no horizon exists, but CTCs are present. This suggests that certain causality-violating events or signals may be observed by external observers. While such a region is permitted by the QIRK metric, its interpretation in the real world raises significant questions and therefore warrants further investigation.
It is also important to note that the black and orange curves intersect at a rotation parameter $a<1$. This implies that, for the QIRK black hole, there exists a maximum rotation parameter $a_{\text{max}}$, which is strictly less than 1. This behavior is different from the Kerr black hole, where extremality is achieved at $a=1$. The specific values of $a_{\text{max}}$ are provided in Table 1.
In this paper, we focus on the cases of $p=4$ and $p=6$ for the following reasons: first, as illustrated in Fig. 3, as the value of $p$ increases, the parameter range for the QIRK black hole expands, and the orange curve representing the avoidance of CTCs approaches the $a$-axis, with $a_{\text{max}}$ gradually approaching 1. This could potentially obscure certain distinctive features of the QIRK black hole. Second, from the aspect of black hole shadow, which we will discuss in Sec. V, a smaller value of $p$ results in greater deviations from the Kerr black hole. Thus, considering smaller values of $p$ better highlights the differences between the QIRK black hole and the Kerr black hole.
Figure 3: Parameter plane $(a,\widetilde{\omega})$ for the QIRK spacetime. The black curve corresponds to $\widetilde{\omega}_{c}$, which separates the black hole spacetimes from the no-horizon spacetimes. The orange curve represents $\widetilde{\omega}_{m}$, which divides spacetimes with closed timelike curves (CTCs) from those without CTCs. The grey region and white region represent black holes and naked singularities without CTCs, respectively. The orange region and green region represent black holes and naked singularities with CTCs, respectively.
Table 1: Maximum rotation parameter $a_{\text{max}}$ and critical rotation parameter $a_{\text{cri}}$ for different values of $p$.
p 4 6 8 10 12 14 16
$a_{\text{max}}$ 0.913141 0.947498 0.964022 0.973442 0.979401 0.983448 0.986342
$a_{\text{cri}}$ 0.873588 0.921738 0.9466 0.961165 0.970453 0.976748 0.981217
III Null geodesics and Photon sphere
When photons move around a black hole, the strong gravitational field causes them to bend. Among the many possible orbits, the most unique is the spherical photon orbit, which defines the boundary between the shaded and the illuminated region. Therefore, studying the motion of photon is crucial for analyzing the structure and observational properties of the black hole shadow.
Before proceeding with the discussion, we would like to clarify the following: we have observed that for many rotating black holes, their null geodesic equations, including the photon sphere, share a similar form to that of the Kerr black hole, with the only difference being a correction to $\Delta$. Considering that, first, we have already discussed the behavior of $\Delta$ for the QIRK black hole in Fig. 1, and second, substituting the specific form of $\Delta$ for the QIRK black hole would make the expressions unnecessarily lengthy and less intuitive, we will not include the explicit form of $\Delta$ in this section. This approach simplifies the generalization to other rotating black holes. Moreover, the analytical expressions derived from the analysis of the photon sphere structure in this section serve as a foundation for intuitively identifying the photon sphere in the next section.
Photon trajectories within the spacetime described by the metric (3) are governed by the geodesic equations, which can be derived from the Hamilton-Jacobi equation 1968PhRv..174.1559C :
$\frac{\partial S}{\partial\tau}=-\frac{1}{2}g^{\alpha\beta}\frac{\partial S}{\partial x^{\alpha}}\frac{\partial S}{\partial x^{\beta}},$ (15)
where $\tau$ is the affine parameter along the geodesics, and $S$ is the Jacobi action. The metric (3) exhibits both time translational and rotational invariance, which leads to conserved quantities along the geodesics: the energy $E=-p_{t}$ and the axial angular momentum $L=p_{\varphi}$, where $p_{\mu}$ represents the photon’s four-momentum. The Petrov-type D nature of the metric (3) guarantees the existence of Carter’s separable constant $\mathcal{K}$, allowing the action to be written in the form:
$S=-Et+L\varphi+S_{r}(r)+S_{\theta}(\theta),$ (16)
where $S_{r}(r)$ and $S_{\theta}(\theta)$ are functions only of $r$ and $\theta$, respectively. The geodesic equations can then be written as the following complete set of first-order differential equations for null geodesics 1968PhRv..174.1559C ; Chandrasekhar_1985 :
$\Sigma\frac{\mathrm{d}t}{\mathrm{d}\lambda}$ $=a(L-aE\sin^{2}\theta)+\frac{r^{2}+a^{2}}{\Delta}\Big{[}(r^{2}+a^{2})E-aL\Big{]}\,,$ (17)
$\Sigma\frac{\mathrm{d}\varphi}{\mathrm{d}\lambda}$ $=\frac{L}{\sin^{2}\theta}-aE+\frac{a}{\Delta}\Big{[}(r^{2}+a^{2})E-aL\Big{]}\,,$ (18)
$\Sigma\frac{\mathrm{d}r}{\mathrm{d}\lambda}$ =±$=\pm\sqrt{R(r)}\,,$ (19)
$\Sigma\frac{\mathrm{d}\theta}{\mathrm{d}\lambda}$ =±$=\pm\sqrt{\Theta(\theta)}\,,$ (20)
where
$R(r)$ $=\Big{[}(r^{2}+a^{2})E-aL\Big{]}^{2}-\Delta\Big{[}\mathcal{K}+(L-aE)^{2}\Big{]}\,,$ (21)
$\Theta(\theta)$ $=\mathcal{K}+\Big{(}a^{2}E^{2}-\frac{L^{2}}{\sin^{2}\theta}\Big{)}\cos^{2}\theta\,.$ (22)
The “+” and “-” signs in Eq. (19) correspond to outgoing and ingoing photons in the radial direction, while in Eq. (20) they represent photons moving towards the south pole $\theta=\pi$ and the north pole $\theta=0$, respectively. The geodesic equations described above govern the propagation of photons in the QIRK spacetime. When studying photon orbits, two distinct impact parameters are typically defined Chandrasekhar_1985 :
$\xi=\frac{L}{E},\quad\eta=\frac{\mathcal{K}}{E^{2}}.$ (23)
These two impact parameters play a crucial role in determining the shape of a photon’s orbit. Based on an analysis of the radial effective potential $\mathcal{V}_{r}(r)$, photon orbits can be classified into three distinct categories: scattered orbits, absorbed orbits, and bounded orbits (also known as spherical orbits or photon sphere). If these spherical orbits are unstable, small perturbations will cause them to be observable, forming an apparent boundary that separates the dark and bright regions in the observer’s sky, ultimately shaping the optical appearance of the black hole. These orbits are determined by the following conditions:
$R(r)=0\,,\quad\quad\frac{\mathrm{d}R(r)}{\mathrm{d}r}=0\,.$ (24)
which can be written exactly with impact parameters $\xi$ and $\eta$
$\frac{R(r)}{E^{2}}$ $=\left(a^{2}-a\xi+r^{2}\right)^{2}-\Delta(r)\left[(\xi-a)^{2}+\eta\right]=0,$ (25)
$\frac{R^{\prime}(r)}{E^{2}}$ $=4r\left(a^{2}-a\xi+r^{2}\right)-\left[(a-\xi)^{2}+\eta\right]\Delta^{\prime}(r)=0.$ (26)
If we are considering the rotating black hole, i.e., $a\neq 0$, the solutions of (25) and (26) are
$\xi$ $=\frac{a^{2}\Delta^{\prime}(r)+r^{2}\Delta^{\prime}(r)-4r\Delta(r)}{a\Delta^{\prime}(r)}\Big{|}_{r=r_{\text{sphere}}}\,,$ (27)
$\eta$ $=\frac{r^{2}\left(16a^{2}\Delta(r)-r^{2}\Delta^{\prime}(r)^{2}+8r\Delta(r)\Delta^{\prime}(r)-16\Delta(r)^{2}\right)}{a^{2}\Delta^{\prime}(r)^{2}}\Big{|}_{r=r_{\text{sphere}}}\,.$ (28)
They are all functions of $r_{\text{sphere}}$, i.e. the radius of the photon sphere. At the same time,
$R^{\prime\prime}(r_{\text{sphere}})$ $=$ $4\left(a^{2}-a\xi+3r^{2}\right)-\left(a^{2}-2a\xi+\eta+\xi^{2}\right)\Delta^{\prime\prime}(r)\Big{|}_{r=r_{\text{sphere}}}\,$ (29)
$=$ $8r\left(\frac{2\Delta(r)\left(\Delta^{\prime}(r)-r\Delta^{\prime\prime}(r)\right)}{\Delta^{\prime}(r)^{2}}+r\right)\Big{|}_{r=r_{\text{sphere}}}\,.$ (30)
If $R^{\prime\prime}(r_{\text{sphere}})>0$, the photon sphere is unstable, while for $R^{\prime\prime}(r_{\text{sphere}})<0$, it is stable.
In the case of a non-rotating black hole ($a=0$), the system exhibits spherical symmetry, allowing us to set $\theta=\pi/2$. In this case, the motion in the $\theta$-direction can be simplified as $\mathcal{K}=0$ (or $\eta=0$). Meanwhile, the impact parameter $\xi$ corresponds to $b$, which is the impact parameter in the spherical symmetric case. Consequently, the equations $R(r)=0$ and $R^{\prime}(r)=0$ reduce to:
$4\Delta(r)-r\Delta^{\prime}(r)=0,$ (31)
$\xi^{2}=\frac{r^{4}}{\Delta(r)}\Big{|}_{r=r_{\text{sphere}}}.$ (32)
In the rotating case, the photon sphere cannot be fully determined by Eqs. (27) and (28) alone. The angular equation of motion (20) must also be considered, specifically requiring $\Theta\geq 0$, where
$\frac{\Theta(\theta)}{E^{2}}=\eta+\left(a^{2}-\frac{\xi^{2}}{\sin^{2}\theta}\right)\cos^{2}\theta\geq 0.$ (33)
Figure 4: Plot of $\Theta(\theta)/E^{2}$ as a function of $\xi$ and $\eta$. There are two types of $\theta$-motion modes. The left panel corresponds to $\xi^{2}\geq a^{2}$, while the right panel corresponds to $\xi^{2}<a^{2}$.
As shown in Fig. 4, the condition $\Theta\geq 0$ leads to two cases:
  • Case 1: When $\xi^{2}\geq a^{2}$ and $\eta\geq 0$, $\Theta(\theta)$ has two zero points, denoted $\theta_{\text{min}}$ and $\theta_{\text{max}}$, such that $\theta_{\text{min}}+\theta_{\text{max}}=\pi$. In this case, the photon oscillates between $\theta_{\text{min}}$ and $\theta_{\text{max}}$, allowing it to cross the equatorial plane. If $\eta=0$, the photon remains confined to the equatorial plane at $\theta=\pi/2$.
  • Case 2: When $\xi^{2}<a^{2}$ and $(a-|\xi|)^{2}+\eta\geq 0$, the behavior differs. Here, $\eta$ may take negative values, altering the $\theta$-motion. For $\eta>0$, the photon’s $\theta$-motion resembles the first case. However, for $\eta<0$, $\Theta(\theta)$ has four distinct zero points: $\theta_{1}<\theta_{2}<\theta_{3}<\theta_{4}$. In this scenario, the photon oscillates within either $(\theta_{1},\theta_{2})$ or $(\theta_{3},\theta_{4})$, and it cannot cross the equatorial plane.
To fully solve for the photon sphere, it is necessary to simultaneously satisfy the three conditions in Eqs. (27), (28), and (33). For photon sphere associated with shadow, these conditions typically reduce to $\eta\geq 0$ in most rotating black holes. In contrast, orbits inside the event horizon may require consideration of cases where $\eta$ is negative.
Once the photon sphere range is determined, by substituting $r_{\text{sphere}}$ into Eqs. (27) and (28), we can compute the impact parameters $\xi$ and $\eta$. The stability of these orbits is then assessed by analyzing $R^{\prime\prime}(r_{\text{sphere}})$, while the permissible range for the photon’s $\theta$-motion is determined by the condition $\Theta\geq 0$, as specified in Eq. (33).
IV Analysis of shadow-related functions
In this section, we will use the analytical expressions for the photon sphere obtained in Sec. III to further investigate the structure of the photon sphere. We will study the asymptotic behavior of the shadow-related functions in a more general case to gain a more intuitive understanding of the photon sphere and its impact on the black hole shadows. Specifically, we will solve for the photon sphere of the QIRK black hole to illustrate these effects.
We consider four functions related to the black hole shadow: $\eta(r)$, $\xi^{2}-a^{2}$, $R^{\prime\prime}(r)$, and $\xi(r)$. The functions $\eta(r)$ and $\xi^{2}-a^{2}$ govern the photon’s mode of motion in the $\theta$-direction, defining the allowed region for photon motion, as discussed in Sec. III. The function $R^{\prime\prime}(r)$ characterizes the stability of the photon sphere, indicating whether a photon sphere at a given radius is stable and observable. Furthermore, $\xi(r)$ directly correlates with the celestial transverse coordinate of the shadow in (42). The monotonicity of $\xi(r)$ within the allowable range of photon sphere radii influences the shape of the shadow. Notably, if $\xi(r)$ exhibits non-monotonic behavior, this may give rise to a ”cuspy-like” shadow structure Cunha:2017eoe .
By examining the expressions for $\xi$ and $\eta$ in (27) and (28), along with the second derivative of $R$ in (30), we can analyze the asymptotic behavior of these functions as $r\rightarrow\infty$, $r\rightarrow r_{+}$, and $r\rightarrow r_{0}$, where $r_{0}$ is the zero point of $\Delta^{\prime}(r)$. Assuming $\Delta(r)\rightarrow r^{n}$ as $r\rightarrow+\infty$, we obtain the following results:
  1. 1.
    As $r\rightarrow\infty$:
    $\xi(r)$ $\rightarrow\frac{n-4}{an}r^{2}\stackrel{{\scriptstyle n=2}}{{\longrightarrow}}-\infty,\quad\eta(r)\rightarrow\frac{-n^{2}+8n-16}{a^{2}n^{2}}r^{4}\stackrel{{\scriptstyle n=2}}{{\longrightarrow}}-\infty,$
    $R^{\prime\prime}(r)$ $\rightarrow\frac{8(4-n)}{n}r^{2}\stackrel{{\scriptstyle n=2}}{{\longrightarrow}}+\infty,\quad\xi^{2}-a^{2}\rightarrow\Big{(}\frac{n-4}{an}\Big{)}^{2}r^{4}-a^{2}\stackrel{{\scriptstyle n=2}}{{\longrightarrow}}+\infty.$
    Here, the asymptotic behavior of the four functions depends on the asymptotic form of $\Delta(r)$ as $r\rightarrow\infty$, particularly for QIRK black holes where $n=2$.
  2. 2.
    In the non-extremal case, $r_{-}\neq r_{+}\neq r_{0}$:
    1. (a)
      As $r\rightarrow r_{+}^{+}$:
      $\xi(r_{+})$ $=\frac{a^{2}+r_{+}^{2}}{a}>0,\quad\quad\quad\eta(r_{+})=-\frac{r_{+}^{4}}{a^{2}}<0,$
      $R^{\prime\prime}(r)$ $=8r_{+}^{2}>0,\quad\quad\quad\xi^{2}-a^{2}=2r_{+}^{2}+\frac{r_{+}^{4}}{a^{2}}>0.$
    2. (b)
      As $r\rightarrow r_{0}^{+}$:
      $\xi(r)$ $\rightarrow\frac{-4r_{0}\Delta(r_{0})}{a\Delta^{\prime}(r_{0})}\rightarrow+\infty,\quad\eta(r)\rightarrow\frac{16r_{0}^{2}\Delta(r_{0})(a^{2}-\Delta(r_{0}))}{a^{2}\Delta^{\prime}(r_{0})^{2}}\rightarrow-\infty,$
      $R^{\prime\prime}(r)$ $\rightarrow-\frac{16r_{0}^{2}\Delta(r_{0})\Delta^{\prime\prime}(r_{0})}{\Delta^{\prime}(r_{0})^{2}}\rightarrow+\infty,\quad\xi^{2}-a^{2}\rightarrow\frac{16r_{0}^{2}}{a^{2}}\Big{(}\frac{\Delta(r_{0})}{\Delta^{\prime}(r_{0})}\Big{)}^{2}-a^{2}\rightarrow+\infty.$
    In this case, the divergence of the four functions at the boundaries $r\rightarrow\infty$ and $r\rightarrow r_{0}$ makes them behave similarly to quadratic functions within this interval. Assuming all four functions are quadratic-like, these asymptotic behaviors allow us to deduce the opening direction of these curves, which provides an intuitive view of the photon sphere structure.
  3. 3.
    In the extremal case, $r_{-}=r_{+}=r_{0}$, when $r\rightarrow r_{0}$ and $\Delta^{\prime\prime}(r_{0})\neq 0$:
    $\xi(r_{0})$ $=\frac{a^{2}+r_{0}^{2}}{a}>0,\quad\eta(r_{0})=r_{0}^{2}\frac{8a^{2}-r_{0}^{2}\Delta^{\prime\prime}(r_{0})}{a^{2}\Delta^{\prime\prime}(r_{0})},$
    $R^{\prime\prime}(r_{0})$ $=0,\quad\quad\quad\xi^{2}-a^{2}=2r_{0}^{2}+\frac{r_{0}^{4}}{a^{2}}>0.$
    Here, the value of $\eta(r_{0})$ depends on the specific black hole model, and its sign determines whether the photon sphere can extend continuously inside the event horizon. This may subsequently affect the analysis of the black hole shadow. Further discussion on this will be provided in Sec. V.
In each case above, the first arrow or equality represents the general scenario, while the following expression is specific to the QIRK black hole. Moreover, by analyzing their monotonicity, we find that:
$R^{\prime\prime}(r)$ $=8r\left(\frac{2\Delta(r)\left(\Delta^{\prime}(r)-r\Delta^{\prime\prime}(r)\right)}{\Delta^{\prime}(r)^{2}}+r\right),$ (34)
$\xi^{\prime}(r)$ $=-\frac{2}{a}\left(\frac{2\Delta(r)\left(\Delta^{\prime}(r)-r\Delta^{\prime\prime}(r)\right)}{\Delta^{\prime}(r)^{2}}+r\right),$ (35)
$\eta^{\prime}(r)$ $=\frac{4r(4a^{2}-4\Delta(r)+r\Delta^{\prime}(r))}{a^{2}\Delta^{\prime}(r)}\left(\frac{2\Delta(r)\left(\Delta^{\prime}(r)-r\Delta^{\prime\prime}(r)\right)}{\Delta^{\prime}(r)^{2}}+r\right).$ (36)
These functions all share the same factor. When this factor equals zero at some $r\neq 0$, both $R^{\prime\prime}(r)$ and $\xi^{\prime}(r)$ will simultaneously vanish. However, $\eta^{\prime}(r)$ cannot be determined due to an additional term. This suggests that the stability of the photon sphere may change along with the monotonicity of $\xi(r)$, as can be observed in the extremal case.
Figure 5: $R^{\prime\prime}(r)$, $\eta(r)$, $\xi(r)$, and $\xi^{2}-a^{2}$ as functions of the photon sphere radius $r_{\text{sphere}}$. The vertical black dashed line represents the horizon. The top two panels show the QIRK BH in the non-extremal case and the Kerr BH with the corresponding rotation parameter $a$. The behavior of the four functions in both panels is highly similar, suggesting that the shadows of these black holes will exhibit comparable characteristics. The bottom two panels depict the extremal QIRK BH and the extremal Kerr BH, both of which display analogous behavior. The behavior of extremal QIRK black holes is discussed in further detail in Sec. V.
In Fig. 5, we have plotted the functions $R^{\prime\prime}(r)$, $\eta(r)$, $\xi(r)$, and $\xi(r)^{2}-a^{2}$ for both the QIRK black hole and the Kerr black hole, which align with the expected asymptotic behavior. Fig. 5 also provides a clear visualization of the photon sphere range and their properties.
In the non-extremal case, the behavior of these four shadow-related functions for the QIRK black hole closely resembles that of the Kerr black hole. Specifically, the range of $r_{\text{sphere}}$ lies outside the horizon, where $R^{\prime\prime}(r)>0$, indicating that they are unstable and directly contribute to the formation of the black hole shadow. Given the similar behavior of these functions in both the QIRK and Kerr black holes, we can infer that the shadow of the QIRK black hole will closely resemble that of the Kerr black hole. This similarity will be further explored in Sec. V.
In the extremal case, however, the functions $\eta(r)$, $\xi^{2}-a^{2}$, $R^{\prime\prime}(r)$, and $\xi(r)$ do not diverge at $r_{0}$ any more. Analyzing the photon region, we observe that $\eta(r_{0})>0$, implying that the region where $\eta(r)\geq 0$, corresponding to the photon region, extends continuously inside the horizon. Moreover, we find $R^{\prime\prime}(r_{0})=0$, indicating a transition in the stability of the photon sphere: while the photon sphere outside the horizon remains unstable, the one inside the horizon is stable. It is important to note that this stable photon sphere does not contribute to the shadow. As illustrated in Fig. 5, this structure closely resembles that of the extremal Kerr black hole.
If we consider the photon sphere located at the horizon, the equations (25) and (26) simplify to:
$\frac{R(r_{0})}{E^{2}}$ $=\left(a^{2}-a\xi+r_{0}^{2}\right)^{2}=0,$ (37)
$\frac{R^{\prime}(r_{0})}{E^{2}}$ $=4r_{0}\left(a^{2}-a\xi+r_{0}^{2}\right)=0.$ (38)
There is no restriction on $\eta$. This means that as long as the photon’s impact parameter $\xi$ satisfies Eq. (37) and Eq. (38), the photon can move in spherical orbits at the horizon.
V Shadow
In general, photons emitted from a light source are deflected by a black hole due to gravitational lensing. These photons can be categorized into three types: scattered, bounded, and absorbed. Scattered photons reach distant observers, while bounded and absorbed photons orbit near the black hole and fall into it, respectively. This creates a dark region on the observer’s plane, known as the black hole shadow. The boundary between the light and dark regions is defined by the orbits of these bounded photons.
For an observer at the position $(r_{obs},\theta_{0})$, in the far exterior region of the black hole, The celestial coordinates $(\alpha,\beta)$ are obtained by projecting the velocity of the photons observed by the observer onto the horizontal and vertical coordinates of the observation plane 1973blho.conf..215B ; Chandrasekhar_1985 , i.e.,
$\alpha=-r_{obs}\frac{p^{\varphi}}{p^{t}},\quad\beta=r_{obs}\frac{p^{\theta}}{p^{t}},$ (39)
where $p^{\mu}$ is the photon four-momentum measured. For an observer sitting in the asymptotically flat region $r_{obs}\rightarrow\infty$, substitute into the geodesic equation, celestial coordinates can be simplified as
$\alpha$ $=$ $\lim_{r_{obs}\to\infty}\Big{(}-r^{2}_{obs}\sin\theta_{0}\frac{\mathrm{d}\varphi}{\mathrm{d}r}\bigg{|}_{\theta=\theta_{0}}\Big{)}=-\frac{\xi}{\sin\theta_{0}}\,,$ (40)
$\beta$ $=$ $\lim_{r_{obs}\to\infty}\Big{(}r^{2}_{obs}\frac{\mathrm{d}\theta}{\mathrm{d}r}\bigg{|}_{\theta=\theta_{0}}\Big{)}=\pm\sqrt{\eta+a^{2}\cos^{2}\theta_{0}-\xi^{2}\cot^{2}\theta_{0}}\,,$ (41)
where $r_{obs}$ denotes the distance between the observer and the black hole, $\theta_{0}$ represents the inclination angle between the observer’s line of sight and the rotational axis of the rotating black hole. When the observer is on the equatorial plane of the black hole with the inclination angle $\theta_{0}=\pi/2$, we obtain
$\alpha$ $=-\xi\,,$ (42)
$\beta$ =±$=\pm\sqrt{\eta}\,.$ (43)
In this situation, the celestial coordinates reduce to the impact parameters of the photons, which are shadow-related functions discussed in Sec. IV.
V.1 Non-Extremal Case
Figure 6: The Quantum Improved Regular Kerr black hole shadows with varying parameters $a$, $p$ and $\widetilde{\omega}$.
In the non-extremal case $(\widetilde{\omega}_{m}<\widetilde{\omega}<\widetilde{\omega}_{c})$, Fig. 6 shows the shadows of QIRK black hole that for fixed rotation $a$ and $p$. As $\widetilde{\omega}$ increases from $\widetilde{\omega}_{m}$ to near the critical value $\widetilde{\omega}_{c}$, we find that the leftmost point of intersection of the shadow with the $\alpha$-axis gradually moves inward, deviating from both the Kerr case and the standard circular shadow. However, the rest of the shadow remains mostly unchanged. Additionally, for fixed $a$ and $\widetilde{\omega}$, increasing $p$ causes the shadows to converge toward those of the Kerr black hole with the same rotation $a$. This suggests that $p=4$ represents the maximum deviation from the Kerr shadow. When $p$ and $\widetilde{\omega}$ are fixed, varying $a$ produces effects similar to those seen in general rotating black holes – larger values of $a$ cause the shadow to deviate more from the standard circular shape.
Overall, in the non-extremal case, smaller values of $p$ and larger values of $\widetilde{\omega}$ lead to greater deviations from the Kerr shadow, particularly where the leftmost point of the shadow intersects the $\alpha$-axis, while the rest of the shadow remains relatively unchanged regardless of $p$ and $\widetilde{\omega}$.
V.2 Extremal Case
In the extremal case $(\widetilde{\omega}=\widetilde{\omega}_{c})$, the inner and outer horizons coincide. For Kerr black holes, achieving extremality requires $a=1$. However, the QIRK black hole can reach the extremal case for any allowable $a$ by setting $\widetilde{\omega}=\widetilde{\omega}_{c}$, as indicated by the black curve in Fig. 3. This extremal case presents notable differences from the non-extremal scenario. Fig. 5 illustrates the shadow-related functions for an extremal QIRK black hole with $p=4$ and $a=0.9$, contrasted with the extremal Kerr black hole. In this extremal configuration, none of the four functions diverge near the horizon.
When considering photon sphere that contribute to the shadow, only cases with $\eta(r)>0$ outside the horizon are relevant, as the photon sphere within the horizon is stable and thus not observable. Consequently, the range of $r_{\text{sphere}}$ should not extend inside the horizon. Notably, however, we observe that $\eta(r_{+})>0$, meaning that when plotting the shadow’s curve with $r_{\text{sphere}}$ as a parameter, the resulting curve will not be closed, since its vertical coordinate $\beta(r_{+})=\sqrt{\eta(r_{+})}>0$. This behavior stems from the unique properties of the photon sphere at the horizon. As discussed in Sec. IV, the photon sphere at the horizon imposes a constraint on $\xi$, while $\eta$ can vary freely, forming a straight line parallel to the $\beta$-axis in the observation plane. This line intersects the open curve, ultimately completing it as a closed loop.
This behavior, however, does not hold across all extremal cases. Fig. 7 shows that for fixed $p$, there exists a critical value $a_{\text{cri}}$, above which the region with $\eta(r)>0$ extends into the horizon. For $a<a_{\text{cri}}$, the scenario resembles that of the non-extremal case, as suggested by our discussion of the extremal case for $\eta(r_{0})$ in Sec. IV. The specific values of $a_{\text{cri}}$ for each $p$ are listed in Table 1.
In Fig. 8, the first panel provides a clear visualization of the discussion in Sec. IV. The dashed lines in all four panels mark the formation of photon sphere at the horizon. Shadows only appear on these dashed lines when $a>a_{\text{cri}}$, not when $a<a_{\text{cri}}$, which aligns with the behavior of the extremal Kerr case. Additionally, the trends in shadow shape and size are consistent with those observed in the non-extremal case: smaller $p$ and larger $a$ result in shadows that deviate further from the standard circular form.
Figure 7: In the extremal case, the relationship between $\eta(r)$ and the horizon with fixed $p$. Different curves correspond to $\eta(r)$ for various rotation parameters $a$, where the vertical lines indicate the positions of the horizon for each corresponding $a$. There exists a critical rotation parameter $a_{\text{cri}}$. For $a>a_{\text{cri}}$, the region where $\eta(r)>0$ extends into the horizon, while for $a<a_{\text{cri}}$, this region remains outside the horizon.
Figure 8: Shadows of the extremal QIRK black hole with varying parameters $a$ and $p$. The solid curves represent the shadow formed by the photon sphere outside the horizon, while the dashed lines correspond to the photon sphere located at the horizon. When $a>a_{\text{cri}}$, both dashed lines and solid curves appear. Conversely, for $a<a_{\text{cri}}$, only solid curves are present.
Overall, the effect of the parameter $p$ and the quantum correction $\widetilde{\omega}$, introduced by the QIRK metric, on black hole shadows follows the same trend for both non-extremal and extremal cases. Specifically, smaller $p$ and larger $\widetilde{\omega}$ within the allowed range $(\widetilde{\omega}_{m}<\widetilde{\omega}\leq\widetilde{\omega}_{c})$ lead to greater deviations from Kerr black holes. However, in terms of the magnitude of this effect on the shadows, when the rotation parameter $a$ is fixed and $p$ and $\widetilde{\omega}$ are varied across their full ranges, the change in the shadows is relatively small compared to many other rotating black holes.
Moreover, the change in the shadow of the QIRK black hole is primarily concentrated at the left intersection of the shadow with the $\alpha$-axis, similar to other rotating black holes based on asymptotic safe gravity Held:2019xde . In other words, even when the quantum effect is significant, its impact on the black hole shadow appears minimized, making it less prominent.
Another noteworthy point is the indistinguishability of extremal QIRK black holes. In the case of extremal Kerr black holes, the strictly straight line on the left side of the shadow makes them distinguishable from their non-extremal counterparts. However, for QIRK black holes, when $a<a_{\text{cri}}$, the shadow of the extremal QIRK black hole does not exhibit a strictly straight line at the leftmost edge, which is similar to the non-extremal case. As a result, it becomes indistinguishable from the non-extremal case. Only when $a>a_{\text{cri}}$ does the extremal QIRK black hole displays a shadow resembling that of the extremal Kerr black hole. Therefore, it is impossible to determine whether a QIRK black hole is in the extremal state based solely on the shape of its shadow.
VI Constraints from EHT Observations
In this section, we utilize observational results from the Event Horizon Telescope (EHT) for the supermassive black holes M87* and Sgr A* to constrain the parameters of the QIRK black hole model. By doing so, we aim to refine the QIRK black hole model, aligning it more closely with astronomical observations. To achieve this, we analyze three shadow observables: the circularity deviation $\Delta C$, the shadow angular diameter $\theta_{d}$, and the Schwarzschild deviation $\delta$, each constrained by EHT data.
The black hole shadow boundary is characterized as a one-dimensional closed curve described by radial and angular coordinates $(R(\varphi),\varphi)$ in a polar coordinate system centered at $(\alpha_{C},\beta_{C})$. The shadow’s average radius $\bar{R}$ is defined as follows johannsen_testing_2010 :
[…]
with
$R(\varphi)=\sqrt{(\alpha-\alpha_{C})^{2}+(\beta-\beta_{C})^{2}},\quad\tan(\varphi)=\frac{\beta-\beta_{C}}{\alpha-\alpha_{C}},$ (45)
where $(\alpha_{C},\beta_{C})$ represent the shadow’s displacement from the black hole center at $(0,0)$. Given the intrinsic axisymmetry, the vertical displacement is zero ($\beta_{C}=0$), and the horizontal displacement $\alpha_{C}$ is
$\alpha_{C}=\frac{\alpha_{r}+\alpha_{l}}{2},$ (46)
where $\alpha_{r}$ and $\alpha_{l}$ denote the abscissae where the shadow intersects the $\alpha$-axis.
The circularity deviation $\Delta C$, measuring the deviation from a perfect circle, is defined as the root-mean-square distance from the average radius johannsen_testing_2010 ; johannsen_photon_2013 ; walia_testing_2022 :
[…]
where $\Delta C=0$ for a perfectly circular shadow. From EHT observations, the circularity deviation for M87* has been constrained to $\Delta C\leq 0.10$ EventHorizonTelescope:2019dse ; EventHorizonTelescope:2019ggy ; EventHorizonTelescope:2019pgp .
The shadow’s angular diameter $\theta_{d}$, defined in terms of the shadow cone’s opening angle, is given by afrin_parameter_2021 :
$\theta_{d}=\frac{2}{D_{LS}}\sqrt{\frac{A}{\pi}},$ (48)
where $A$ is the shadow area, expressed as abdujabbarov_coordinate-independent_2015 :
[…]
Additionally, the observable $\delta$, quantifying the deviation between the observed angular diameter and that of a Schwarzschild black hole, is defined as EventHorizonTelescope:2022xqj ; EventHorizonTelescope:2022urf :
$\delta=\frac{\theta_{d}}{\theta_{d,Sch}}-1,$ (50)
where $\theta_{d,Sch}$ is the angular diameter of a Schwarzschild black hole’s shadow.
In the following, we assume that M87* and Sgr A* are QIRK black holes and aim to constrain the model parameters based on astronomical data. By analyzing these observables, we will exclude parameter ranges incompatible with the data, using density maps based on Fig. 3. While uncertainties in the EHT observations of M87* and Sgr A* introduce some degree of imprecision, we can still estimate the approximate effect of the parameters introduced by the QIRK black holes in the observations.
In Sec. V, we analyze the black hole shadow assuming an inclination angle of $\theta_{0}=90^{\circ}$, which maximizes deviation from circularity and enhances the effects of the black hole’s properties. However, various inclination angles can be considered for observational constraints. For instance, EHT estimates the inclination angle of M87* to be approximately $163^{\circ}$, based on the relativistic jet orientation CraigWalker:2018vam . Since the shadow is symmetric about the $\alpha$-axis, $\theta_{0}=163^{\circ}$ is equivalent to $\theta_{0}=17^{\circ}$ in our analysis. For Sgr A*, while the inclination angle is not definitively known, values below $52^{\circ}$ are generally favored. Inclinations of $5^{\circ}$ Kuang:2022ojj , $50^{\circ}$ walia_testing_2022 , and $90^{\circ}$ Ghosh:2022kit have been employed in previous analyses. The circularity deviation $\Delta C$ increases with the inclination angle $\theta_{0}$, so we can obtain an upper bound on the parameters by choosing an inclination of $90^{\circ}$. However, the shadow area decreases as the inclination angle increases, and the angular diameter $\theta_{d}$ and Schwarzschild deviation $\delta$ are proportional to the area. Therefore, smaller inclination angles can provide tighter constraints when considering $\theta_{d}$ and $\delta$. After taking these factors into account, we select $\theta_{0}=17^{\circ}$ and $90^{\circ}$ for M87*, and $50^{\circ}$ for Sgr A*.
VI.1 Constrains from M87*
The first image of the supermassive black hole M87* revealed an asymmetric bright ring caused by strong gravitational lensing and relativistic beaming, with a central dark region identified as the black hole shadow EventHorizonTelescope:2019dse ; EventHorizonTelescope:2019ggy ; EventHorizonTelescope:2019pgp . Given the distance of M87* from Earth, $D_{\text{M87*}}=16.8$ Mpc, and its estimated mass $M_{\text{M87*}}=(6.5\pm 0.7)\times 10^{9}M_{\odot}$, constraints can be placed on the emission region. The circularity deviation is constrained to $\Delta C\leq 0.10$, with an angular diameter of $\theta_{d,\text{M87*}}=(42\pm 3)\mu\text{as}$ and a Schwarzschild deviation of $\delta_{\text{M87*}}=-0.01\pm 0.17$ within a $1\sigma$ confidence interval.
Figures 9, 10, and 11 display the density plots of $\Delta C$, $\theta_{d}$, and $\delta$ under varying inclinations and parameters. As shown in Fig. 9, $\Delta C$ remains below 0.10 for $\theta_{0}=17^{\circ}$, providing no substantial constraints. However, for $\theta_{0}=90^{\circ}$, some parameters with $\Delta C>0.10$ are excluded. From Fig. 10, we can observe that the area of the shadow $A$ increases as the inclination angle $\theta_{0}$ decreases. Therefore, $\theta_{0}=17^{\circ}$ provides stronger constraints if we exclude $\theta_{d}<39\mu\text{as}$. For $\theta_{d}$, the exclusion criteria are $\theta_{d}>45\mu\text{as}$ or $\theta_{d}<39\mu\text{as}$ within the $1\sigma$ interval. Given the modest deviation of the QIRK black hole from the Kerr black hole, we do not extend the analysis to the $2\sigma$ interval. Finally, Fig. 11 shows that all values of $\delta$ lie within the $1\sigma$ interval, meaning no significant exclusion can be made based on $\delta$ alone.
Figure 9: The shadow circularity deviation observable $\Delta C$ for different inclination angles $\theta_{0}$ and parameter $p$, as a function of $(a,\widetilde{\omega})$. The dashed blue curve indicates $\Delta C=0.10$, where the region to the right of this curve is excluded based on the observed circularity deviation of the M87* black hole as reported by the EHT, $\Delta C\leq 0.10$.
Figure 10: The shadow angular diameter $\theta_{d}$ for QIRK black holes, with different inclination angles $\theta_{0}$ and parameter $p$, as a function $(a,\widetilde{\omega})$. For M87*, EHT observations report an angular diameter of $\theta_{d}=42\pm 3\,\mu\text{as}$. The dashed blue curve corresponds to $\theta_{d}=39\,\mu\text{as}$, and the region to the right of this curve exceeds the $1\sigma$ range of the observational result, thus being excluded.
Figure 11: Schwarzschild shadow angular diameter deviation $\delta$ as a function of $(a,\widetilde{\omega})$. The dashed blue curve represents $\delta=-0.01$.
VI.2 Constrains from Sgr A*
In the published results for Sgr A*, the EHT provides measurements for both the angular diameter of the emission ring, $\theta_{d,\text{Sgr~{}A*}}=(51.8\pm 2.3)\mu\text{as}$, and the angular diameter of the black hole shadow, $\theta_{sh,\text{Sgr~{}A*}}=(48.7\pm 7)\mu\text{as}$. Given the broad range of the shadow angular diameter, we will use an average value of $\theta_{d}\in(46.9,50)\mu\text{as}$, as determined by the three independent imaging algorithms employed by the EHT walia_testing_2022 . For uniformity, we will denote all shadow angular diameters simply as $\theta_{d}$. The Schwarzschild shadow deviation is given by $\delta_{\text{Sgr~{}A*}}=-0.08^{+0.09}_{-0.09}$ (VLTI) and $\delta_{\text{Sgr~{}A*}}=-0.04^{+0.09}_{-0.10}$ (Keck) at the $1\sigma$ confidence level. In the following analysis, we adopt the mass of Sgr A* as $M_{\text{Sgr~{}A*}}=4.0^{+1.1}_{-0.6}\times 10^{6}M_{\odot}$ and its distance from Earth as $D_{\text{Sgr~{}A*}}=8.15\pm 0.15$ kpc EventHorizonTelescope:2022wkp ; EventHorizonTelescope:2022xqj ; EventHorizonTelescope:2022urf ; EventHorizonTelescope:2022wok ; EventHorizonTelescope:2022apq ; EventHorizonTelescope:2022exc .
For Sgr A*, only the angular diameter $\theta_{d}$ and deviation $\delta$ are considered, as the EHT did not provide a bound on the circularity deviation $\Delta C$. As with M87*, Fig. 12 allows us to exclude parameters where $\theta_{d}>50\mu\text{as}$. Additionally, Fig. 13 plots $\delta=-0.04$ (Keck), but this does not yield significant constraints.
Fig. 14 provides a unified representation of the constraints from all observables. For M87*, the curve $\theta_{d}=39\,\mu\text{as},\,\theta_{0}=17^{\circ}$ excludes a significant number of parameters, particularly those to the right of it. In contrast, for Sgr A*, parameters to the left of $\theta_{d}=50\,\mu\text{as},\,\theta_{0}=50^{\circ}$ are excluded. In Fig. 14, the blue curves represent the most stringent constraint on the parameter range, while the gray regions indicate the ultimately allowed parameter range.
After synthesizing the results from M87* and Sgr A*, we obtain a band-like region between $\theta_{d}=39\,\mu\text{as}$, $\theta_{0}=17^{\circ}$ and $\theta_{d}=50\,\mu\text{as}$, $\theta_{0}=50^{\circ}$. From this, we observe that for the Kerr black hole, there are upper and lower bounds on the rotation parameter $a$. However, the introduction of quantum corrections alters these bounds. In general, the larger the correction $\widetilde{\omega}$, the smaller both the upper and lower bounds become. Once the correction $\widetilde{\omega}$ exceeds a certain value, the lower bound of $a$ can even reach zero, which is significantly different from the Kerr black hole. Moreover, when $a=0$, $\widetilde{\omega}$ is constrained to a region near the $\widetilde{\omega}_{c}$, with most of its values being excluded. The precise intersection points between each curve and the $\widetilde{\omega}_{c}$, $\widetilde{\omega}_{m}$, or $\widetilde{\omega}$-axis are provided in Table 2 and Table 3, corresponding to M87* and Sgr A*, respectively.
Figure 12: Shadow angular diameter $\theta_{d}$ for QIRK black holes with different values of parameter $p$, plotted as a function of $(a,\widetilde{\omega})$. The dashed blue curve represents $\theta_{d}=50\mu\text{as}$, corresponding to the Sgr A* black hole shadow bounds. QIRK black holes in the region to the right of the blue curve produce shadows consistent with the observed size of the Sgr A* shadow.
Figure 13: Schwarzschild shadow angular diameter deviation $\delta$ as a function of $(a,\widetilde{\omega})$. The dashed blue curve corresponds to $\delta=-0.04$ (Keck).
Figure 14: The observational bounds for M87* and Sgr A* with different $p$. The black and green curves represent the extremal case $(\widetilde{\omega}_{c})$, separated by $a_{\text{cri}}$ (the critical value for the shadow in the extremal case). The orange curve corresponds to $\widetilde{\omega}_{m}$, with the intersection of the two curves indicating $a_{\text{max}}$, the maximum value of $a$ for which both the black hole and the absence of CTCs are satisfied, as discussed in Sec. II. In the first panels, parameters to the right are excluded by the blue curve $\theta_{d}=39\,\mu\text{as},\,\theta_{0}=17^{\circ}$, the red dashed curve $\theta_{d}=39\,\mu\text{as},\,\theta_{0}=17^{\circ}$, and the magenta dashed curve $\Delta C=0.10,\,\theta_{0}=90^{\circ}$. In the second panel, parameters to the right are excluded by the red dashed curve $\theta_{d}=39\,\mu\text{as},\,\theta_{0}=90^{\circ}$ and the blue curve $\Delta C=0.10,\,\theta_{0}=90^{\circ}$. In contrast, in the lower two panels, parameters to the left are excluded by the blue curve $\theta_{d}=50\,\mu\text{as},\,\theta_{0}=50^{\circ}$. In summary, the final parameter constraints are represented by the blue curves, which provide the most restrictive limits on the parameters. Meanwhile, the gray regions represent the final allowed parameter range.
Table 2: Intersection points of observational constraints with the curves $\widetilde{\omega}_{c}$, $\widetilde{\omega}_{m}$, or the $\widetilde{\omega}$-axis for M87*.
$p=4$ $p=6$
$\Delta C=0.1$ $\theta_{d}=39\mu\text{as}$ $\theta_{d}=39\mu\text{as}$ $\delta=-0.01$ $\delta=-0.01$ $\Delta C=0.1$ $\theta_{d}=39\mu\text{as}$ $\delta=-0.01$
$\theta_{0}=90^{\circ}$ $\theta_{0}=90^{\circ}$ $\theta_{0}=17^{\circ}$ $\theta_{0}=90^{\circ}$ $\theta_{0}=17^{\circ}$ $\theta_{0}=90^{\circ}$ $\theta_{0}=90^{\circ}$ $\theta_{0}=90^{\circ}$
$\widetilde{\omega}_{c}$ $(0.537,2.361)$ $(0.591,2.034)$ $(0.472,2.748)$ $(0.352,3.386)$ $(0.277,3.717)$ $(0.537,2.361)$ $(0.675,3.374)$ $(0.504,6.544)$
$\widetilde{\omega}_{m}$ $(0.633,0.075)$ $(0.717,0.115)$ $(0.553,0.047)$ $(0.568,0.051)$ $(0.426,0.018)$ $(0.633,0.075)$ $(0.724,0.058)$ $(0.571,0.020)$
Table 3: Intersection points of observational constraints with the curves $\widetilde{\omega}_{c}$, $\widetilde{\omega}_{m}$, or the $\widetilde{\omega}$-axis for Sgr A*.
$p=4,\theta_{0}=50^{\circ}$ $p=6,\theta_{0}=50^{\circ}$
$\theta_{d}=39\mu\text{as}$ $\delta=-0.01$ $\theta_{d}=39\mu\text{as}$ $\delta=-0.01$
$\widetilde{\omega}_{c}$ / $\widetilde{\omega}$axis $(0,3.882)$ $(0.855,0.513)$ $(0.304,10.093)$ $(0.875,0.695)$
$\widetilde{\omega}_{m}$ $(0.374,0.011)$ $(0.874,0.226)$ $(0.374,0.003)$ $(0.886,0.141)$
VII Conclusion
In this paper, building on the work of Ref. chen2024quantumimprovedregularkerr , we conduct a more detailed study of the QIRK black hole. Due to the difficulty of analytic discussions, we employed numerical methods and, after comprehensively considering the three conditions of resolving singularities, ensuring the existence of horizons, and avoiding closed timelike curves, we determined the parameter range allowed for the QIRK black hole. We found that for a given $p$, the quantum correction $\widetilde{\omega}$ has both upper and lower bounds ($\widetilde{\omega}_{c}$ and $\widetilde{\omega}_{m}$). This also indicates that, under these three conditions, the quantum effects cannot be too large or too small. Additionally, the rotational parameter $a$ has a maximum value less than 1, which differs significantly from the Kerr black hole, indicating that the QIRK black hole does not exhibit extremely high rotation. Moreover, we also discovered that the QIRK metric can describe a compact object with closed timelike curves but no horizon. The causality-violating signals from this region could potentially be observable by external observers, making this area worthy of further investigation.
As a strong field test, the black hole shadow can reveal both universal and unique features of different black hole models. To study the shadow of the QIRK black hole, we conducted a comprehensive analysis of the photon sphere structure in a more general context and provided the analytical expressions for the photon sphere. Through the shadow-related function, we constructed an intuitive image of the photon sphere structure, establishing a specific correspondence between the photon sphere and the black hole shadow. We also specifically considered the photon sphere of the QIRK black hole in both non-extremal and extremal cases, and compared it to the Kerr black hole. From the plots of the shadow-related function for the non-extremal case, we observe that their behaviors are quite similar, indicating that their shadows differ only in size. In particular, in the extremal case, the photon sphere at the horizon exhibits distinctive properties, which are reflected in the shadow.
The introduction of the quantum correction parameter $\widetilde{\omega}$ extends the extremal black hole condition from $a=1$ in Kerr black holes to a curve $\widetilde{\omega}=\widetilde{\omega}_{c}$ in the QIRK black hole. This gives the extremal QIRK black hole a richer structure, manifested in the fact that when $a<a_{\text{cri}}$, the shadow of the extremal QIRK black hole resembles that of a non-extremal black hole, making them indistinguishable. Only when $a>a_{\text{cri}}$ does the shadow of the extremal QIRK black hole resembles that of the extremal Kerr black hole. Regardless of whether the black hole is in an extremal state, the trends in shadow changes with $p$ and $\widetilde{\omega}$ remain the same, differing only in magnitude. The overall changes remain small, indicating that the quantum effects introduced by the QIRK model do not produce significant alterations in the black hole shadow.
When considering the recent observational results of M87* and Sgr A* from the EHT to further investigate the QIRK black hole as a potential candidate for astrophysical observations, we find that, due to the relatively small impact of quantum effects on the shadow, even when considering all allowed parameters, the three observables reflecting the shadow size—$\Delta C$, $\theta_{d}$, and $\delta$— still exhibit variations within a relatively small range. To obtain stronger constraints, we consider different inclination angles. For the case of $p=4$, the permissible parameters lie within the region to the left of the blue curve in Fig. 14a, and to the right of the blue curve in Fig. 14c. Both regions are shaded in grey in the figure. The final result is the intersection of these two grey regions, forming a band-like area. A similar region is formed for $p=6$, following the same pattern. Thus, based on the current results, the QIRK black hole remains a viable candidate for observational testing.
In this paper, we only discussed the cases of $p=4$ and $p=6$. This choice was made for two reasons. First, a larger $p$ increases the degree of the polynomial, making the analysis more complicated. Second, as $p$ increases, the shadow of the QIRK black hole gradually converges to that of the Kerr black hole, making them indistinguishable. Moreover, current observational data would not be sufficient to constrain the parameters in such cases. Therefore, we focused on the two smallest values of $p$ to highlight the distinctive features of this black hole model and to obtain stronger constraints.
Whether the QIRK black hole can serve as a viable candidate for a real astrophysical black hole will ultimately depend on more precise and abundant observational data. Future observations will be crucial in either verifying or ruling out the QIRK black hole as a realistic black hole model.
Acknowledgement
We would like to thank Nobuyoshi Ohta for his useful discussion on the QIRK black holes. This work was supported in part by the National Natural Science Foundation of China with grants No.12075232 and No.12247103.